OpenStudy (anonymous):

what is the slope and y intercept of a line that contains the point (5,-1) and has the same intercept as x-3y=6?

OpenStudy (anonymous):

Find x and y intercept of x-3y=6

OpenStudy (anonymous):

then use Slope intercept form

OpenStudy (anonymous):

would x= 6 and y=-2?

OpenStudy (nubeer):

yes it would be that @Judith5

OpenStudy (anonymous):

how would i plug them in the slope intercept form?

OpenStudy (hba):

The slope intercept form is y-y1=m(x-x1)

OpenStudy (anonymous):

would i put the x= 6 and y=-2 to that equation or the points given in the question?

OpenStudy (hba):

No (x1,y1)=(6,-2)

OpenStudy (nubeer):

i dont think so @hba 6 and -2 were the x and y intercepts.. not a point.

OpenStudy (hba):

Oh I did not see . (x1,y1)=(5,-1) Sorry about that.

OpenStudy (hba):

You can find the slope also which you have to by the formula m=-coefficient of x/co-efficient of y

OpenStudy (anonymous):

so i would put (6-2)=(5-1)?

OpenStudy (hba):

No put the point (x1,y1)=(5,-1) in y-y1=m(x-x1)

OpenStudy (hba):

I gave you the formula for the slope also so first find slope. Then use the equation.

OpenStudy (hba):

@Judith5 Got it ?

OpenStudy (anonymous):

yup, thanks

OpenStudy (hba):

Your'e welcome.